Solve the following
y+3 = 2x
x^2 +2xy = -1
Can you please help me out? Thank you

Answer by mananth(16946)

y+3 = 2x
y=2x-3
x^2 +2xy = -1
x^2+2xy+1=0
(x+y)^2=0
substitute y
(x+2x-3)^2=0
(3x-3)^2=0
9(x-1)^2=0
(x-1)^2=0
x-1=0
x=1
y+3=2x
x=1
y+3=2
y=-1
x=1,y=-1